# First-Principles Calculations on Elasto-optical Properties of RRTe3

### v1

@@ I Introduction
Rare-earth tritellurides ($R\mathrm{Te}_{3}$, where $R=\mathrm{La}-\mathrm{Yb}$) are a family of quasi-two-dimensional materials with tunable lattice and electronic properties[1, 2]. Each compound crystallizes in an orthorhombic $Cmcm$ structure (space group $D_{2h}$), consisting of nearly square-planar Te layers separated by corrugated rare-earth–Te slabs along the $b$-axis[3, 1]. The weak van der Waals coupling between layers and the small in-plane anisotropy between the $a$ and $c$ axes make these materials highly responsive to external strain, providing a versatile platform for studying strain-tunable electronic order[4, 5].
@@ III.4 Symmetry analysis
Unconventional charge-density-wave (CDW) states in RRTe3 have been proposed to host additional symmetry-breaking beyond the orthorhombic $D_{2h}$ (space group Cmcm) lattice — for example, ferroaxial or monoclinic order that reduces the point group[10, 9, 7]. Such hidden order may not produce a large change in conventional transport, but it imposes new selection rules on optical and elasto-optical tensors; this makes piezo-optical measurements a sensitive probe of additional symmetry lowering associated with an unconventional CDW transition[13, 22].

### v2

@@ I Introduction
Rare-earth tritellurides ($R\mathrm{Te}_{3}$, where $R=\mathrm{La}-\mathrm{Yb}$) are a family of quasi-two-dimensional materials with tunable lattice and electronic properties[1, 2]. Each compound crystallizes in an orthorhombic $Cmcm$ structure (space group $D_{2h}$), consisting of nearly square-planar Te layers separated by corrugated rare-earth–Te slabs along the $b$-axis[3, 1]. The weak van der Waals coupling between layers and the small in-plane anisotropy between the $a$ and $c$ axes make these materials highly responsive to external strain, providing a versatile platform for studying strain-tunable electronic order[4, 5, 6].
@@ III.4 Symmetry analysis
Unconventional charge-density-wave (CDW) states in RRTe3 have been proposed to host additional symmetry-breaking beyond the orthorhombic $D_{2h}$ (space group Cmcm) lattice — for example, ferroaxial or monoclinic order that reduces the point group[11, 10, 8]. Such hidden order may not produce a large change in conventional transport, but it imposes new selection rules on optical and elasto-optical tensors; this makes piezo-optical measurements a sensitive probe of additional symmetry lowering associated with an unconventional CDW transition[14, 23, 24].